### v1

# ContrastScore: Towards Higher Quality, Less Biased, More Efficient Evaluation Metrics with Contrastive Evaluation

Abstract
Evaluating the quality of generated text automatically remains a significant challenge. Conventional reference-based metrics have been shown to exhibit relatively weak correlation with human evaluations. Recent research advocates the use of large language models (LLMs) as source-based metrics for natural language generation (NLG) assessment. While promising, LLM-based metrics, particularly those using smaller models, still fall short in aligning with human judgments. In this work, we introduce ContrastScore, a contrastive evaluation metric designed to enable higher-quality, less biased, and more efficient assessment of generated text. We evaluate ContrastScore on two NLG tasks: machine translation and summarization. Experimental results show that ContrastScore consistently achieves stronger correlation with human judgments than both single-model and ensemble-based baselines. Notably, ContrastScore based on Qwen 3B and 0.5B even outperforms Qwen 7B, despite having only half as many parameters, demonstrating its efficiency. Furthermore, it effectively mitigates common evaluation biases such as length and likelihood preferences, resulting in more robust automatic evaluation.
1 Introduction
Evaluating the quality of automatically generated text remains a fundamental challenge in natural language processing (NLP) and, in some cases, is nearly as difficult as generating the text itself. Traditional evaluation methods primarily rely on reference-based metrics such as BLEU, ROUGE, and METEOR  Papineni et al. (2002); Lin (2004); Banerjee and Lavie (2005); Shen et al. (2022); Pan et al. (2024); Kalyan (2024), which are inadequate as they compare generated text to human-written references based on surface-level lexical overlap, often failing to capture semantic adequacy and fluency. Embedding-based metrics like BERTScore Zhang et al. (2019) improve upon lexical approaches by leveraging contextualized representations for better semantic alignment. However, they remain sensitive to domain shifts, depend on high-quality references, and exhibit relatively weak correlations with human judgments Zhao et al. (2023).
To address these limitations, recent research has shifted towards source-based evaluation methods, leveraging large language models (LLMs). Metrics such as BARTScore (Yuan et al., 2021) and GPTScore (Fu et al., 2024) assess generation quality using a single model’s probability distribution, while prompt-based LLM evaluators process structured prompts to evaluate text based on predefined criteria Que et al. (2024) or conduct comparative judgments without an explicit rubric Liusie et al. (2023). Although these methods show promise, their effectiveness is inherently constrained by the underlying model’s capacity. Larger models generally perform better but are computationally expensive and incur high API costs (Larionov and Eger, 2024). Conversely, smaller models often have reduced capacity, resulting in unreliable evaluations. Furthermore, existing LLM-based metrics are susceptible to biases (Deutsch et al., 2022; Sun et al., 2022; Liu et al., 2024), including length bias and over-reliance on likelihood, which may not always align with human evaluation criteria. These challenges highlight the need for novel evaluation methods that are both efficient and reliable, balancing model capacity, accessibility, and evaluation robustness (Chen and Eger, 2023; Zhao et al., 2024).
Introduced by Li et al. (2023), Contrastive Decoding enhances both diversity and factual accuracy by leveraging the disparity between a stronger expert model and a weaker amateur model, prioritizing tokens with the largest probability gap. Beyond decoding, contrastive principles have also been applied to model architectures, such as the Self-Contrast Mixture-of-Experts (SCMoE), where different expert pathways within a model act as internal contrastive filters to enhance reasoning Shi et al. (2024). These approaches highlight the potential of contrastive mechanisms in improving model outputs by using weaker models to generate contrastive signals, enabling stronger models to refine their predictions and achieve a balance between fluency and diversity.
Inspired by contrastive principles, we introduce ContrastScore, an evaluation metric that leverages structured disagreement between two models of differing capacities. Unlike conventional LLM-based evaluation methods that rely solely on a single model’s likelihood estimates, ContrastScore incorporates a weaker auxiliary model as a contrastive signal, dynamically adjusting probability scores to improve alignment with human judgements. Specifically, ContrastScore is built on a discrepancy-based probability formulation that measures the absolute difference between the probabilities assigned by two models. By utilizing discrepancies between a stronger (expert) model and a weaker (amateur) model, it generates more calibrated and robust evaluation scores.
We conduct extensive experiments on two widely employed text generation tasks, namely, machine translation and summarization, to evaluate the effectiveness of ContrastScore. Our evaluation compares ContrastScore against established metrics and various baseline models, including single-model and ensemble-based approaches. Experimental results show that ContrastScore achieves a higher correlation with human judgments, outperforming both single-model and ensemble-based methods. Notably, ContrastScore using smaller models (Qwen 3B, Qwen 0.5B) even surpasses Qwen7B with 5.2% of improvement in summarization despite having only half the parameters, demonstrating its efficiency. Furthermore, ContrastScore based on smaller models substantially enhances evaluation speed, providing at least a 1.5-fold increase in processing speed compared to a single larger model across both the Qwen and Llama families. Additionally, it effectively mitigates length and likelihood biases, enhancing robustness in automatic evaluation. By incorporating contrastive principles into evaluation, ContrastScore paves the way for a new paradigm of more robust and efficient text evaluation.
To summarize, the contributions of our paper are four-fold:
  • We propose a simple yet highly effective additive-based formulation for contrastive evaluation, leveraging structured model discrepancies to produce more calibrated and reliable evaluation scores.
  • We conduct extensive experiments across multiple generation tasks, diverse datasets, and various model families to rigorously assess the effectiveness of our approach.
  • Contrastive evaluation strongly correlates with human judgments, outperforming existing source-based metrics while effectively addressing likelihood and length biases, which are prevalent in automatic evaluation.
  • ContrastScore achieves significantly faster evaluation compared to larger single-model approaches, improving inference speed by at least 1.5 times while using only half the parameters.
2 Related Work
LLM-Based Evaluation.
With the rise of LLMs, evaluation has shifted towards using LLMs as direct evaluators of text generation quality. Among LLM-based approaches, BARTScore [^1] The original BARTScore used BART as a language model, but this can be replaced by any more recent LLM such as GPT4 (Fu et al., 2024). Yuan et al. (2021) has emerged as a widely used and reliable metric due to its flexibility and strong correlation with human judgments. It leverages a pretrained model to compute log-likelihood scores for generated text, making it a strong indicator of text quality.
Beyond BARTScore, another prominent evaluation paradigm involves using LLMs as evaluators via prompting. This approach, commonly referred to as LLM-as-a-Judge, directly queries LLMs with carefully designed prompts to assess the quality of generated text. One notable framework following this paradigm is G-Eval Liu et al. (2023), which systematically prompts LLMs to evaluate text across multiple quality attributes (e.g. coherence and consistency), allowing for a more structured and adaptable assessment. Similarly, ChatGPT-based evaluation Gao et al. (2023) explores the use of ChatGPT as a human-like summarization evaluator, demonstrating its ability to approximate human judgments. These methods leverage the implicit knowledge and reasoning abilities of LLMs to provide fine-grained evaluations, making them well-suited for tasks where traditional metrics fall short. ChatEval Chan et al. (2024) introduces a multi-LLM evaluation framework based on a multi-agent debate system. Unlike single-model evaluators, ChatEval orchestrates multiple LLM-based agents, each assuming different roles and perspectives.
Shortcoming of LLM-based Evaluation.
Prompting-based evaluation techniques suffer from prompt sensitivity, where small changes in the input prompt can lead to significantly different results. This sensitivity raises concerns about result consistency, as even slight rewordings or rephrasings of the prompt can alter the model’s response Leiter and Eger (2024). Furthermore, these approaches rely on the capabilities of the specific LLM used, meaning that while they may perform well when evaluated with state-of-the-art models, the reliability of their assessments deteriorates when used with weaker models.
More broadly, LLM-based evaluation methods inherit the biases and limitations of the underlying model. One key issue is likelihood bias, where LLMs tend to favor sentences with higher likelihood scores based on their internal language modeling prior, regardless of the task’s requirements or the actual quality of the generated text Ohi et al. (2024). This bias leads to inflated scores for fluent but incorrect or hallucinated outputs, undermining the evaluation’s reliability. Furthermore, LLMs may exhibit biases inherited from their training data, leading to systematic errors in assessing factuality, coherence, or stylistic appropriateness.
Contrastive Decoding.
Contrastive methods have recently gained attention in text generation, where the differences between models of varying capacities are leveraged to improve output quality Su et al. (2022). Rather than relying on ensemble methods or majority voting, contrastive decoding explicitly uses a weaker model as a contrastive signal to refine or steer the output of a stronger model. This approach has demonstrated success in mitigating common generation issues such as repetition, overconfidence, and incoherence. Li et al. (2023) utilizes a large ‘expert’ model and a smaller ‘amateur’ model. Instead of directly following the expert’s probabilities, contrastive decoding selects tokens that maximize the log-probability difference between the two models. This results in text that is more fluent, diverse, and factually accurate. Moreover, contrastive principles have been extended to model architecture optimization. Shi et al. (2024) propose the Self-Contrast Mixture-of-Experts (SCMoE) framework, where different expert pathways within the same model act as internal contrastive filters, which improves reasoning capabilities and overall model performance, eliminating the need for external weaker models. Note that contrastive decoding is also related to language model arithmetic in which arithmetic combinations of LLMs are considered for adjusted generation (Dekoninck et al., 2024). This includes subtraction, which leads to output favored by one LLM but disfavored by the other, which could be leveraged in adversarial settings (Zhang and Eger, 2024).
Inspired by these advancements, as described in §3.3, we extend contrastive principles beyond generation to the domain of evaluation. Instead of refining model outputs, we propose ContrastScore—a novel metric with a different formulation from prior contrastive decoding work. By leveraging a discrepancy-based probability framework that quantifies the absolute difference between the probabilities assigned by two models, our approach generates contrastive signals that enable more calibrated and reliable evaluation scores.
3 Methodology
The primary goal of this study is to introduce a contrastive evaluation framework, ContrastScore. Current generative metrics for text evaluation are predominantly based on a single language model. However, their performance is highly dependent on the capabilities of the underlying model, leading to performance disparities, where metrics based on smaller or medium-sized models typically underperform those using larger models. Additionally, these single-model approaches are inherently affected by the biases of the underlying model such as likelihood bias, where the model may over-prefer fluent but less accurate outputs, or pretrained biases that can influence evaluation outcomes in unintended ways.
To address this, we introduce ContrastScore, a method that leverages the structured disagreement between two models of different capacities to produce a bias-aware, dynamically calibrated probability score for evaluation. Rather than relying solely on the probability of a single model, ContrastScore incorporates an auxiliary model as a contrastive signal, allowing for a more refined probability adjustment that improves alignment with human judgments.
3.1 Generative Evaluation
Our evaluation framework builds on probability-based text assessment methods, such as BARTScore, which estimate quality through log-likelihood computations under a pretrained generative model. Given a generated hypothesis h, its quality is evaluated as:
$\text{Score}(\textbf{h}\mid d,\mathcal{S})=\sum_{t=1}^{m}\log P(h_{t}\mid\textbf{h}_{<t},\mathcal{S},\theta),$ (1)
where $\mathcal{S}$ represents the supplementary text, which may consist of the source text s (in a source-based setting) or the reference output r (in a reference-based setting). $\theta$ denotes the model parameters, and $d$ is the description of the task instruction.
3.2 Contrastive Decoding
Assume an expert and amateur model assigning probabilities
$p_{\text{EXP}}^{t}=p(h_{t}\mid h_{<t},\mathcal{S},\theta_{\text{EXP}})$ (2)
$p_{\text{AMA}}^{t}=p(h_{t}\mid h_{<t},\mathcal{S},\theta_{\text{AMA}})$ (3)
to the next token $h_{t}$, where $p_{\text{EXP}}^{t}$ and $p_{\text{AMA}}^{t}$ represent the probabilities of the expert and amateur models, respectively, for $h_{t}$ and where $\theta_{\text{EXP}}$ and $\theta_{\text{AMA}}$ refer to the parameters of the expert and amateur model. Li et al. (2023) propose the contrastive decoding objective
[…]
Here, $V_{\text{head}}(h_{<t})$ selects the top most likely tokens $x_{t}$ under the expert model given history $h_{<t}$. During next token prediction, CD-Score effectively only considers the most likely tokens of the expert model as candidates, and then assigns modified probabilities $\log{\frac{p_{\text{EXP}}^{t}}{p_{\text{AMA}}^{t}}}$ to them. The intuition behind CD-Score is to consider the most likely expert tokens as continuation for generation but then choose those tokens among them which the amateur might not favor (e.g., repetitive tokens).
Unfortunately, CD-Score is unsuitable for evaluation because every generated sequence h containing a token $h_{t}$ not in $V_{\text{head}}(h_{<t})$ would receive a score of $\infty$- ∞, essentially being ruled out as a candidate, even though it should have been considered as part of the sequence under evaluation. As an alternative, one might consider the division score $\log{\frac{p_{\text{EXP}}^{t}}{p_{\text{AMA}}^{t}}}$ itself as an objective for evaluation. However, as pointed out by Li et al. (2023), this approach suffers from penalizing many standard text sequences as unlikely, for which both expert and amateur would hold the same probabilities. For example, if both models assign a probability of 0.9 or 0.1 to a token, the score remains the same, making it unable to distinguish between confidently correct and incorrect predictions. As a result, this approach fails to highlight important distinctions between tokens that should matter in evaluation. The division score also becomes instable especially when the amateur probability is close to zero, distorting the evaluation score.
3.3 ContrastScore
To address this issue for evaluation, we propose a subtraction-based contrastive formulation with a scaling factor $\gamma\in[0,1]$ that downweights the amateur model. In this way, we naturally retain all tokens in the input sequence for evaluation, as well as a property that favors generations preferred by the expert model, similar to CD-Score. Importantly, the subtraction-based formulation ensures that the scores of individual tokens remain comparable across positions—unlike the division-based approach, where small variations in the amateur model’s likelihood can lead to disproportionately large and unstable scores. By using a small $\gamma$, we control the influence of the amateur model’s probability and mitigate such instability, yielding more reliable evaluation signals.
Empirically, we choose a formula for contrastive evaluation that leverages the absolute value distance between the expert and the scaled down amateur:
$\text{ContrastScore}=\sum_{t=1}^{m}\log\left(\left|p_{\text{EXP}}^{t}-\gamma p_{\text{AMA}}^{t}\right|\right)$ (5)
At each time step $t$, ContrastScore rewards tokens $h_{t}$ for which expert and (downweighted) amateur generation probabilities are maximally distinct (as the absolute value is a distance function between two distributions). This would ensure that ContrastScore leverages the strengths of the expert model but removes the limitations of the amateur model; in practice, it might for example lean toward generations that are less repetitive, which the amateur model tends to favor. Note, however, that there are probability ranges where the scaled down amateur would be preferred over the expert — for example, when $p^{t}_{\text{EXP}}=0$, then our formula would still assign the scaled probability of the amateur to $h_{t}$. With $\gamma=1$, ContrastScore would be indifferent between the expert and the amateur but indiscriminately favor disagreement between them, an undesirable behavior similar to the division score above.
4 Experimental Setup
We evaluate the performance and robustness of the proposed ContrastScore on two tasks: machine translation and summarization, using models from the Llama and Qwen families to ensure generalizability. To benchmark ContrastScore, we compare it against single-model evaluators as well as an ensemble setting, where the probabilities of the expert and amateur models are averaged per token. The ensemble probability at token $t$, denoted as $p_{\text{Ens}}^{t}$, is computed as: $p_{\text{Ens}}^{t}=\frac{1}{2}\left(p_{\text{EXP}}^{t}+p_{\text{AMA}}^{t}\right)$.
4.1 Datasets and Tasks
For machine translation experiments, we utilize the WMT23 Metrics Shared Task dataset Freitag et al. (2023), focusing on three language pairs: English-German (EN-DE), Chinese-English (ZH-EN), and Hebrew-English (HE-EN). These pairs were selected to cover a diverse range of translation scenarios, including high-resource (EN-DE), typologically distant (ZH-EN), and low-resource (HE-EN) settings, ensuring a comprehensive evaluation of our approach. For summarization, we use the SummEval dataset Fabbri et al. (2021), which consists of summaries generated from CNN/DailyMail articles.
4.2 Meta-Evaluation
We assess the evaluation metrics in terms of quality, biases, and efficiency.
Correlation with Human Scores (Quality).  The effectiveness of an evaluator is measured using the Pearson correlation between its scores and human scores, where a higher correlation indicates stronger alignment with human judgments. For translation, Pearson correlation assesses the evaluator’s ability to capture differences between candidate translations based on the source texts, comparing them to human judgment scores derived from fine-grained MQM annotations (Freitag et al., 2021). For summarization, we compute Pearson correlation with human judgments across four quality aspects: coherence, fluency, consistency, and relevance.
Bias Evaluation.  We measure biases in two dimensions: likelihood bias and length bias.
Likelihood Bias: Recent studies have shown that LLMs tend to favor sentences they deem more likely, often assigning higher evaluation scores and performing better on such sentences, regardless of the specific task or evaluation criteria Ohi et al. (2024); McCoy et al. (2024a, b). Ohi et al. (2024) find that LLM-based evaluators systematically overrate high-likelihood sentences and underrate low-likelihood ones compared to human scores. To quantify this likelihood bias, they propose BiasScore, defined as:
$\text{BiasScore}=\rho(\textit{LS},\textit{US})$ (6)
where $\rho$ is the Spearman correlation coefficient, LS (Likelihood Score) represents the probability $P$ assigned by the LLM, and US (Unfairness Score) captures the discrepancy between evaluator scores and human scores. Likelihood bias measures the extent to which an evaluation metric is influenced by the model’s inherent probability estimates. A lower BiasScore indicates that the metric is less dependent on the model’s likelihood biases.
Length Bias: It is observed that summarization evaluation metrics tend to favor longer summaries, particularly in source-based settings Sun et al. (2019); Liu et al. (2024). Liu et al. (2024) find that LLM-based evaluators exhibit this length bias, often assigning higher scores to longer texts regardless of quality. Similarly, we quantify this tendency using the Spearman correlation between evaluator scores and the length of the generated text, where a lower correlation magnitude indicates a weaker length bias, suggesting a more content-focused evaluation.
Efficiency.  An evaluator’s efficiency is measured by the number of samples processed per second, with a higher processing speed indicating greater efficiency.
Source-based Reference-based
Llama EN-DE ZH-EN HE-EN Avg EN-DE ZH-EN HE-EN Avg
1B 0.530 0.518 0.320 0.456 0.608 0.499 0.456 0.521
Single 3B 0.578 0.526 0.293 0.466 0.628 0.498 0.466 0.531
8B 0.632 0.574 0.491 0.566 0.636 0.498 0.474 0.536
(3B,1B) 0.568 0.538 0.306 0.471 0.629 0.499 0.454 0.528
Ensemble (8B,1B) 0.605 0.575 0.462 0.547 0.639 0.505 0.461 0.535
(8B,3B) 0.619 0.574 0.453 0.549 0.645 0.505 0.466 0.539
(3B,1B) 0.599 0.562 0.284 0.482 0.643 0.532 0.463 0.546
ContrastScore (8B,1B) 0.641 0.590 0.482 0.571 0.653 0.541 0.484 0.559
(8B,3B) 0.639 0.595 0.482 0.572 0.652 0.542 0.486 0.560
(a) Correlation of Llama-based evaluators with human scores.
Source-based Reference-based
Qwen EN-DE ZH-EN HE-EN Avg EN-DE ZH-EN HE-EN Avg
0.5B 0.487 0.557 0.347 0.463 0.577 0.509 0.469 0.518
Single 3B 0.573 0.594 0.415 0.527 0.599 0.475 0.477 0.517
7B 0.600 0.574 0.445 0.540 0.608 0.480 0.478 0.522
(3B,0.5B) 0.548 0.599 0.421 0.523 0.610 0.503 0.473 0.528
Ensemble (7B,0.5B) 0.567 0.598 0.446 0.537 0.616 0.505 0.476 0.532
(7B,3B) 0.599 0.608 0.456 0.554 0.624 0.502 0.482 0.536
(3B,0.5B) 0.590 0.628 0.431 0.550 0.627 0.531 0.480 0.546
ContrastScore (7B,0.5B) 0.611 0.628 0.465 0.568 0.632 0.535 0.489 0.552
(7B,3B) 0.605 0.629 0.466 0.567 0.629 0.535 0.490 0.551
(b) Correlation of Qwen-based evaluators with human scores .
Table 1: Correlation of Llama and Qwen-based evaluators with human scores on the WMT23 dataset. Correlation values that are highest within each group (i.e. Single, Ensemble and ContrastScore) are underlined. Additionally, the highest correlation for each language pair is boldfaced.
4.3 Hyperparameters and Environment
To ensure robust generalization, we use models from the same family but of different sizes, where a larger model serves as the expert and a smaller model as the amateur. Specifically, we use LLama3.2-Instruct (1B, 3B) and LLama3.1-Instruct (8B), as the LLama 3.2 version does not include an 8B model. Additionally, we evaluate Qwen2.5-Instruct (0.5B, 3B, and 7B). We also attempted to explore ContrastScore using models from different families; however, this proved problematic due to differences in model tokenization. In all our experiments [^2] We set the decoding temperatures for both the expert and amateur models, denoted as $\tau_{1}$ and $\tau_{2}$, respectively. Following Li et al. (2023), we set $\tau_{2}=0.5$, while $\tau_{1}$ is set to 1.5., the scaling factor for the amateur model, $\gamma$, is set to 0.1. This setting is based on our pilot investigation, which found that $\gamma=0.1$ gives the best overall performance empirically for the ZH-EN language pair of the machine translation task, see Appendix B, Figure 2. All experiments were run on a GPU cluster under same conditions and node configuration to ensure fair comparisons. Each node of the cluster has 4$\times$×H100 GPUs.
5 Results
We assess the effectiveness of ContrastScore across multiple aspects, including its correlation with human scores, bias mitigation, and computational efficiency, in comparison to single-model and ensemble methods.
Llama Coherence Consistency Fluency Relevance Avg.
1B 0.417 0.549 0.542 0.344 0.463
Single 3B 0.388 0.575 0.530 0.338 0.458
8B 0.368 0.588 0.539 0.332 0.457
(3B,1B) 0.427 0.564 0.542 0.344 0.469
Ensemble (8B,1B) 0.429 0.581 0.556 0.357 0.481
(8B,3B) 0.414 0.587 0.544 0.351 0.474
(3B,1B) 0.456 0.554 0.551 0.334 0.474
ContrastScore (8B,1B) 0.461 0.584 0.575 0.368 0.497
(8B,3B) 0.462 0.587 0.575 0.369 0.498
(a) Correlation of Llama-based evaluators with human scores
Qwen Coherence Consistency Fluency Relevance Avg.
0.5B 0.386 0.536 0.520 0.315 0.439
Single 3B 0.370 0.544 0.492 0.319 0.431
7B 0.371 0.560 0.510 0.327 0.442
(3B,0.5B) 0.394 0.568 0.536 0.331 0.457
Ensemble (7B,0.5B) 0.386 0.577 0.538 0.333 0.459
(7B,3B) 0.378 0.584 0.521 0.330 0.453
(3B,0.5B) 0.397 0.594 0.546 0.323 0.465
ContrastScore (7B,0.5B) 0.376 0.608 0.554 0.322 0.465
(7B,3B) 0.384 0.606 0.557 0.334 0.470
(b) Correlation of Qwen2.5-based evaluators with human scores
Table 2: Correlation of Llama and Qwen-based evaluators with human scores on the SummEval dataset. Correlation values that are relatively highest within each group (i.e. Single, Ensemble and ContrastScore) are underlined. Additionally, the highest correlation for each evaluation dimension is boldfaced.
5.1 Correlation with Human Scores
Machine Translation.  Table 1 presents the Pearson correlation between various evaluators and human scores for both source-based and reference-based settings, following Qin et al. (2023). We additionally present a visual comparison in Fig. 4 in Appendix C. The results show that ContrastScore consistently outperforms both single-model and ensemble approaches for both settings. For example, ContrastScore based on Qwen(7B, 0.5B) improves correlation by 5.2% over the single Qwen 7B model and by 5.8% over the ensemble of Qwen(7B, 0.5B) in the source-based setting. It reveals that ensembling does not always improve performance. While ensemble methods are generally expected to refine predictions, they can sometimes reduce correlation with human scores. For instance, ensembling Llama (8B, 3B) in the source-based setting leads to a 3.0% decrease in correlation compared to the single Llama 8B model. Similarly, the ensemble of Llama (3B, 1B) in reference-based setting results in a marginal 0.6% reduction compared to the single Llama 3B model. Additionally, we observe that ContrastScore’s effectiveness depends on the reliability of the amateur model. When the amateur model fails to provide a meaningful correction signal, ContrastScore may not yield improvements. This is evident in the HE-EN pair for Llama models in the source-based setting, where the correlation scores for both the 1B and 3B models (0.32 and 0.29, respectively) are too low to serve as reliable correction signals. Consequently, ContrastScore does not surpass the single 8B model’s correlation in this case.
Beyond these cases where the expert model underperforms relative to the amateur model or when the amateur model is too weak to provide a meaningful correction signal, our results demonstrate that ContrastScore can achieve correlation scores comparable to, or even exceeding, those of the largest model—using only two smaller models. This has significant efficiency implications, as evaluation metrics based on LLMs are used frequently throughout the model development lifecycle, and inference costs scale with model size. Since larger models are computationally expensive, being able to obtain reliable evaluation with smaller models presents a practical advantage. This efficiency gain is particularly valuable for scalable and cost-effective evaluation pipelines. We further explore this trade-off in §5.3.
Summarization.   Table 2 presents the Pearson correlation between different evaluators and human scores for summarization. We additionally present a visual comparison in Fig. 5 in Appendix C. Similar to prior work Liu et al. (2023); Fu et al. (2024), we do not include reference-based evaluation for summarization, as source-based evaluation has been shown to be more robust. Reference-based metrics often struggle with the high variability of reference summaries, leading to weaker correlations with human judgments. In contrast, source-based approaches better assess factual consistency and relevance based on more informative sources, making them more reliable for evaluating summarization quality.
Our results indicate that ContrastScore provides best results overall. For instance, ContrastScore with Llama(8B, 3B) improves correlation by 9.0% over the single Llama 8B, whereas ensembling Llama(8B, 3B) only provides a 3.7% gain. This further supports the effectiveness of ContrastScore as a superior alternative to ensembling. Beyond overall performance, ContrastScore consistently surpasses single models while utilizing significantly fewer parameters. Specifically, ContrastScore with Qwen(3B, 0.5B) outperforms the single Qwen 7B by 5.2% , and ContrastScore with Llama(3B, 1B) exceeds the single Llama 8B by 3.7%. These results reinforce the observation that ContrastScore can efficiently enhance model evaluation by leveraging smaller models.
Besides improving overall correlation with human scores, ContrastScore demonstrates notable gains across key summarization dimensions, particularly coherence, fluency, and relevance. When applied to Llama(8B, 3B), it enhances coherence by 25.5%, fluency by 6.7%, and relevance by 11.1% compared to the single Llama 8B. However, its impact on consistency differs across model families. While it consistently improves consistency scores in Qwen-based models, its effectiveness in Llama-based models is less predictable. This suggests that the way ContrastScore interacts with different evaluation dimensions may depend on model-specific characteristics.
Discussion.  Overall, ContrastScore consistently outperforms both single-model and ensemble approaches across diverse settings, demonstrating its robustness in aligning with human judgments. While it enhances evaluation performance across tasks, it is particularly effective in summarization, substantially improving coherence, fluency, and relevance. In machine translation, it also provides notable gains, especially for high-resource language pairs. Moreover, ContrastScore maintains strong evaluation performance while leveraging smaller models, making it a computationally efficient alternative to single-model methods.
5.2 Bias Analysis
Likelihood Bias. Table 3 presents the likelihood bias scores across different evaluators. The results clearly demonstrate that ContrastScore substantially mitigates likelihood bias compared to both single-model and ensemble approaches. Specifically, ContrastScore with Llama (8B, 3B) reduces the likelihood bias by 37.0% and 61.1% compared to single Llama 8B in summarization and machine translation task, respectively. Similarly, ContrastScore with Qwen (7B, 0.5B) results in a reduction of 25.6% and 34.8% compared to single Qwen 7B model for the same tasks.
Llama Summ WMT Qwen Summ WMT
Single 1B 0.348 0.212 0.5B 0.376 0.252
3B 0.385 0.245 3B 0.392 0.451
8B 0.381 0.352 7B 0.398 0.463
Ensemble (3B,1B) 0.308 0.123 (3B,0.5B) 0.379 0.327
(8B,1B) 0.326 0.152 (7B,0.5B) 0.393 0.294
(8B,3B) 0.359 0.229 (7B,3B) 0.369 0.381
Contrast (3B,1B) 0.183 0.026 (3B,0.5B) 0.236 0.314
(8B,1B) 0.262 0.134 (7B,0.5B) 0.296 0.302
(8B,3B) 0.240 0.137 (7B,3B) 0.329 0.287
Table 3: Likelihood bias scores for machine translation and summarization tasks across Qwen and Llama model families. The relatively highest bias value within each group is underlined. The lowest overall bias score is boldfaced.
Llama Summ WMT Qwen Summ WMT
Single 1B 0.218 -0.692 0.5B 0.289 -0.686
3B 0.289 -0.529 3B 0.391 -0.525
8B 0.344 -0.500 7B 0.408 -0.460
Ensemble (3B,1B) 0.212 -0.633 (3B,0.5B) 0.349 -0.590
(8B,1B) 0.274 -0.628 (7B,0.5B) 0.373 -0.562
(8B,3B) 0.301 -0.564 (7B,3B) 0.403 -0.505
Contrast (3B,1B) 0.086 -0.677 (3B,0.5B) 0.253 -0.599
(8B,1B) 0.206 -0.632 (7B,0.5B) 0.323 -0.565
(8B,3B) 0.201 -0.632 (7B,3B) 0.329 -0.579
Table 4: Length bias scores for machine translation and summarization tasks. The relatively highest bias value within each group is underlined. The lowest overall bias score is boldfaced.
Length Bias.  Table 4 shows the length bias scores for different models across summarization and machine translation tasks. We observe that ContrastScore considerably reduces length bias in summarization. Notably, ContrastScore with Llama (3B,1B) achieves a substantial reduction of 70.2%, lowering the bias score from 0.289 to 0.086. Similarly, for Qwen (3B,0.5B), ContrastScore reduces the bias score from 0.349 to 0.253, demonstrating its effectiveness in mitigating length bias in summarization. While the impact of length bias in machine translation is less well understood and has received comparatively limited attention, we include length bias scores in machine translation task for completeness. As expected, we do not observe a clear or consistent trend in the machine translation setting.
5.3 Efficiency Analysis
Llama Summ WMT Qwen Summ WMT
Single 1B 63.29 339.04 0.5B 141.96 552.68
3B 27.08 153.25 3B 27.69 194.22
8B 26.43 72.17 7B 26.43 99.84
Ensemble (3B,1B) 39.51 110.63 (3B,0.5B) 44.94 154.5
(8B,1B) 23.81 66.8 (7B,0.5B) 26.11 93.92
(8B,3B) 19.81 53.95 (7B,3B) 20.35 71.04
Contrast (3B,1B) 39.12 109.31 (3B,0.5B) 44.06 153.99
(8B,1B) 23.33 66.11 (7B,0.5B) 25.71 93.41
(8B,3B) 19.05 53.45 (7B,3B) 19.61 70.97
Table 5: Processing Speed for Machine Translation and summarization evaluation tasks. Measured in Samples Per Second with batch size of 16 on H100 GPUs in bfloat16 precision.
Figure 1: Performance and efficiency of ContrastScore with smaller models compared to single larger model on summarization task.
Table 5 reports processing speeds for both machine translation (WMT) and summarization (Summ) evaluation tasks, measured in samples per second with a batch size 16, running in bfloat16 precision. In the LLaMa family, the single large model LLaMa 8B processes 26.43 samples/s on Summ and 72.17 samples/s on WMT, whereas ContrastScore using two smaller models (LLaMa 3B, 1B) achieves 39.12 samples/s on Summ and 109.31 samples/s on WMT — about 1.5$\times$ faster in both cases. Similarly, in the Qwen family, the single large model Qwen 7B shows 26.43 samples/s on Summ and 99.84 samples/s on WMT, while ContrastScore with Qwen (3B, 0.5B) processes 44.06 samples/s on Summ and 153.99 samples/s on WMT (approximately 1.7$\times$ faster on Summ and 1.5$\times$ faster on WMT). Moreover, as shown in Tables 1(b) and 2, the ContrastScore method with two smaller models (for both model families) also produces higher correlation with human judgments compared to using a single larger model. Figure 1 shows that employing two smaller models in a ContrastScore framework yields both better efficiency and improved evaluation quality.
6 Further Analysis
Original CE-Score COH CON FL RE AVG
(3B,1B) 0.405 0.441 0.432 0.3373 0.404
Llama (8B,1B) 0.388 0.471 0.446 0.329 0.408
(8B,3B) 0.418 0.475 0.476 0.347 0.429
(3B,0.5B) 0.448 0.474 0.463 0.340 0.431
Qwen (7B,0.5B) 0.429 0.487 0.473 0.332 0.430
(7B,3B) 0.431 0.485 0.487 0.336 0.435
Table 6: Pearson correlation between scores based on the original ratio-based Contrastive Decoding formulation (Li et al., 2023) and human judgments, where the abbreviations COH, CON, FL, and RE stand for coherence, consistency, fluency, and relevance, respectively.
Source Text {CJK*}UTF8gbsn产品防伪标识在哪里
Tokens: Where is the anti - fe iting logo $\bar{\textbf{Log(P)}}$ Rank
Hypothesis 1 Human Rank: 1 Expert 0.2471 0.7305 0.9922 0.02881 0.9805 0.7969 1.000 0.000457 -0.717 2
Amateur 2.265e-06 0.5625 0.9922 0.000335 1.000 1.000 1.000 0.06592 -1.319 1
Contrast 0.2471 0.6758 0.8945 0.02881 0.8789 0.6953 0.8984 0.006134 -0.605 1
Tokens: Where is the anti - fe iting product
Hypothesis 2 Human Rank:2 Expert 0.2471 0.7305 0.9922 0.02881 0.9805 0.7969 1.000 0.002808 -0.618 1
Amateur 2.265e-06 0.5625 0.9922 0.000335 1.000 1.000 1.000 5.841e-05 -1.701 2
Contrast 0.2471 0.6758 0.8945 0.02881 0.8789 0.6953 0.8984 0.002808 -0.647 2
Tokens: What Is The National Debt Limit
Hypothesis 3 Human Rank:3 Expert 0.005951 0.000168 0.2910 4.268e-05 0.001602 0.004944 -2.668 3
Amateur 0.000572 6.845e-08 0.1543 6.482e-07 3.123e-05 0.000140 -4.294 3
Contrast 0.005951 0.000168 0.2754 4.268e-05 0.001602 0.004944 -2.672 3
Table 7: Case Study: Comparison of ContrastScore with a single expert model (Qwen 3B) and an amateur model (Qwen 0.5B) for Chinese-to-English (ZH-EN) machine translation. $\bar{\textbf{Log(P)}}$ denotes mean log-probability across all target tokens.
Weighted Ensemble.   We investigate whether the benefits of ContrastScore could be achieved through a simpler method based on probability combination. Specifically, we test a weighted ensemble approach that linearly combines the probabilities from both larger and smaller models.
$p_{\text{Ens}}^{t}=\gamma\cdot p_{\text{EXP}}^{t}+(1-\gamma)\cdot p_{\text{AMA}}^{t}$ (7)
where $\gamma\in[0,1]$ is the weight factor that controls the contribution of each model. When $\gamma=0.5$, this formulation simplifies to the standard ensemble method, which averages probabilities as reported in our main results. To further analyze the impact of $\gamma$, we conduct a parameter sweep on the summarization task using Llama3.2 (3B, 1B), varying $\gamma$ from 0 to 1 with a step size of 0.05. The results, presented in Figure 3 in Appendix C, show that the weighted ensemble achieves its highest performance at $\gamma=0.55$, with an average correlation of 0.4694. However, $\gamma=0.5$ yields nearly identical results, suggesting that a simple averaging method is nearly as effective as the best-weighted combination. More importantly, ContrastScore (0.474) outperforms the optimal setting of weighted ensembling (0.469) in correlation, highlighting its advantage over probability ensembling approaches.
Ratio-based Contrastive Decoding Formulation. To demonstrate that ContrastScore’s discrepancy-based probability formulation is better suited for evaluation tasks than ratio-based contrast decoding formulation in Li et al. (2023), we test Eq. (4) on the summarization evaluation task.
Table 6 presents the Pearson correlation results between the ratio-based formulation in Li et al. (2023) and human scores for both the Llama and Qwen model families. In all cases, the ratio-based scores are inferior to single-model baselines. This can be attributed to the division formulation, where the score approaches 1 when both the expert and amateur models assign either high or low probabilities, thereby reducing the distinguishability for evaluation.
Case Study
To provide a qualitative analysis, Table 7 presents a case study comparing ContrastScore with a single expert model (Qwen 3B) and an amateur model (Qwen 0.5B) for Chinese-to-English (ZH-EN) machine translation. It can be observed that the expert model misranked Hypothesis 1 as the second-best translation, whereas human evaluators ranked it as the best (Rank 1). This misranking was largely influenced by the expert model assigning a very low probability (0.000457) to “logo”, a key token in the translation. In contrast, the amateur model assigned a significantly higher probability (0.06592) to “logo”, demonstrating greater confidence in its relevance. ContrastScore leveraged this probability discrepancy to adjust the probability of “logo”, thereby increasing the overall score of Hypothesis 1 and correctly ranking it as the best translation. This case illustrates how ContrastScore mitigates expert model underestimation of critical tokens, leading to improved evaluation alignment with human judgment.
In the case of Hypothesis 3, all methods (i.e., expert, amateur, and ContrastScore) correctly ranked it as the worst translation (Rank 3). This consistency occurred because both the expert and amateur models assigned low probabilities across all tokens, indicating poor translation quality. Unlike in Hypothesis 1, where significant probability discrepancies required adjustment, ContrastScore made minimal modifications since the expert and amateur models were already in agreement in their evaluation. This demonstrates that ContrastScore does not introduce unnecessary changes when the models agree, ensuring that it only refines scores when meaningful corrections are needed.
7 Conclusion
In this paper, we introduce ContrastScore, an evaluation metric that leverages structured disagreement between two language models of differing capacities. Unlike conventional LLM-based evaluation methods that rely solely on a single model’s likelihood estimates, ContrastScore incorporates a weaker auxiliary model as a contrastive signal to adjust probability scores for better alignment with human judgments. Extensive experiments demonstrate that ContrastScore consistently achieves stronger correlation with human evaluations than both single-model and ensemble-based baselines. Furthermore, ContrastScore is efficient and can effectively mitigate common evaluation biases, such as preferences for length and likelihood, resulting in more robust evaluation.
Appendix A Prompts
The prompts for the summarization and machine translation tasks in source-based setting are presented in Table 8. For the reference-based setting, the prompts are "To rephrase it" and "That is to say".
Tasks Prompts
machine translation Translate the following sentence to {target language}:
[…]
Table 8: Prompts for tasks description
Appendix B The impact of hyperparameter gamma
We study how sensitive our method is to $\gamma$ in Figure 2.
We test ZH-EN language pair of machine translation task, based on Llama family. Figure 2 shows that $\gamma\in[0.08,0.2]$ leads to good performance, robust in different model size settings. Furthermore, $\gamma=0.1$ produces the best performance, ensuring a small, controlled refinement.
Figure 2: Exploration of the impacts of $\gamma$. Testing correlation between evaluator score and human score in ZH-EN language pair of machine translation task.
Appendix C Results
Figure 3: Exploration of weighted ensemble parameter $\gamma$ for Llama-3.2 (3B, 1B) on summarization. Best quality occurs at $\gamma=0.55$ (star), but remains below ContrastScore using the same models. Horizontal lines show individual model performance.
Figure 4: Correlation of various evaluation scores with human score on WMT dataset
Figure 5: Correlation of various evaluation scores with human score on SummEval dataset