### v2

# ContrastScore: Towards Higher Quality, Less Biased, More Efficient Evaluation Metrics with Contrastive Evaluation

Abstract
Recent advances in automatic evaluation of natural language generation have increasingly relied on large language models as general-purpose metrics. While effective, these approaches often require high-capacity models, which introduce substantial computational costs, and remain susceptible to known evaluation pathologies, such as over-reliance on likelihood. We introduce ContrastScore, a contrastive evaluation paradigm that builds on the widely used BARTScore formulation by comparing token-level probabilities between a stronger and a weaker model. Instead of relying on single-model likelihoods or prompt-based judgments, ContrastScore captures disagreement between models to better reflect confidence and uncertainty in generation quality. Empirical results on summarization and machine translation benchmarks show that ContrastScore, instantiated with paired moderate-scale models across both Qwen and LLaMA families, consistently outperforms larger alternatives, such as Qwen 7B and LLaMA 8B, in correlation with human ratings. In addition to improving evaluation quality, ContrastScore significantly reduces susceptibility to likelihood bias, offering a more robust and cost-effective alternative to larger LLM-based evaluation methods.  [^1] Source code is available at https://github.com/sandywangxiao/ContrastScore.
ContrastScore: Towards Higher Quality, Less Biased, More Efficient Evaluation Metrics with Contrastive Evaluation
Xiao Wang1, Daniil Larionov2, Siwei Wu1, Yiqi Liu1, Steffen Eger3, Nafise Sadat Moosavi4, Chenghua Lin1 1The University of Manchester, 2University of Mannheim, 3University of Technology Nuremberg, 4The University of Sheffield {xiao.wang-26,siwei.wu-2,yiqi.liu-6}@postgrad.manchester.ac.uk daniil.larionov@uni-mannheim.de, steffen.eger@utn.de n.s.moosavi@sheffield.ac.uk, chenghua.lin@manchester.ac.uk
1 Introduction
Evaluating the quality of automatically generated text remains a fundamental challenge in natural language processing (NLP) and, in some cases, is nearly as difficult as generating the text itself. Traditional evaluation methods primarily rely on reference-based metrics such as BLEU, ROUGE, and METEOR  Papineni et al. (2002); Lin (2004); Banerjee and Lavie (2005); Shen et al. (2023); Pan et al. (2024); Kalyan (2024), which are inadequate as they compare generated text to human-written references based on surface-level lexical overlap, often failing to capture semantic adequacy and fluency. Embedding-based metrics like BERTScore Zhang et al. (2020) improve upon lexical approaches by leveraging contextualized representations for better semantic alignment. However, they remain sensitive to domain shifts, depend on high-quality references, and exhibit relatively weak correlations with human judgments Zhao et al. (2023). To address these limitations, recent research has shifted towards source-based evaluation methods, leveraging large language models (LLMs). Metrics such as BARTScore (Yuan et al., 2021) and GPTScore (Fu et al., 2024) assess generation quality using a single model’s probability distribution, while prompt-based LLM evaluators process structured prompts to evaluate text based on predefined criteria Que et al. (2024) or conduct comparative judgments without an explicit rubric Liusie et al. (2024). Although these methods show promise, their effectiveness is inherently constrained by the underlying model’s capacity. Larger models generally perform better but are computationally expensive and incur high API costs (Larionov and Eger, 2024). Conversely, smaller models often have reduced capacity, resulting in unreliable evaluations. Furthermore, existing LLM-based metrics are susceptible to biases (Deutsch et al., 2022; Sun et al., 2022; Lu et al., 2023; Liu et al., 2024; Hong et al., 2025), including over-reliance on likelihood, which may not align with human evaluation criteria. These challenges highlight the need for novel evaluation methods that are both efficient and reliable, balancing model capacity, accessibility, and evaluation robustness (Chen and Eger, 2023; Zhao et al., 2024).
Introduced by Li et al. (2023), Contrastive Decoding enhances both diversity and factual accuracy by leveraging the disparity between a stronger expert model and a weaker amateur model, prioritizing tokens with the largest probability gap. Beyond decoding, contrastive principles have also been applied to model architectures, such as the Self-Contrast Mixture-of-Experts (SCMoE), where different expert pathways within a model act as internal contrastive filters to enhance reasoning Shi et al. (2024). These approaches highlight the potential of contrastive mechanisms in improving model outputs by using weaker models to generate contrastive signals, enabling stronger models to refine their predictions and achieve a balance between fluency and diversity.
Inspired by contrastive principles, we introduce ContrastScore, an evaluation metric that leverages structured disagreement between two models of differing capacities. Unlike conventional LLM-based evaluation methods that rely solely on a single model’s likelihood estimates, ContrastScore incorporates a weaker auxiliary model as a contrastive signal, dynamically adjusting probability scores to improve alignment with human judgements. Specifically, ContrastScore is built on a discrepancy-based probability formulation that measures the absolute difference between the probabilities assigned by two models. By utilizing discrepancies between a stronger (expert) model and a weaker (amateur) model, it generates more calibrated and robust evaluation scores. We conduct extensive experiments on two widely employed text generation tasks, namely, machine translation and summarization, to evaluate the effectiveness of ContrastScore. Our evaluation compares ContrastScore against established metrics and various baseline models, including single-model and ensemble-based approaches. Experimental results show that ContrastScore achieves a higher correlation with human judgments, outperforming both single-model and ensemble-based methods. Notably, ContrastScore using smaller models (Qwen 3B, Qwen 0.5B) even surpasses Qwen7B with 7.0% of improvement in summarization despite having only half the parameters, demonstrating its efficiency. Furthermore, ContrastScore based on smaller models substantially enhances evaluation speed, providing at least a 1.5-fold increase in processing speed compared to a single larger model across both the Qwen and LLaMA families. Additionally, it effectively mitigates the likelihood biase, enhancing robustness in automatic evaluation. By incorporating contrastive principles into evaluation, ContrastScore paves the way for a new paradigm of more robust and efficient text evaluation.
The contributions of our paper are four-fold:
  • We propose a simple yet highly effective difference-based formulation for contrastive evaluation, leveraging structured model discrepancies to produce more calibrated and reliable evaluation scores.
  • We conduct extensive experiments across multiple generation tasks, diverse datasets, and various model families to rigorously assess the effectiveness of our approach.
  • Contrastive evaluation strongly correlates with human judgments, outperforming single-model and ensemble methods while effectively addressing likelihood bias, which are prevalent in automatic evaluation.
  • ContrastScore achieves significantly faster evaluation compared to larger single-model approaches, improving inference speed by at least 1.5 times while using only half the parameters, while maintaining comparable or even superior performance.
2 Related Work
Automatic Evaluation of Text Generation.
Automatic evaluation metrics for text generation can broadly be categorized into task-specific and general-purpose approaches. Early task-specific metrics such as BLEU Papineni et al. (2002), ROUGE Lin (2004), and CHRF Popović (2015) rely on surface-level n-gram overlap with a reference. While originally designed for tasks like machine translation and summarization, these metrics often fail to reflect actual semantic similarity due to the variability of valid natural language expressions. To overcome these limitations, embedding-based metrics, such as BERTScore Zhang et al. (2020) and MoverScore Zhao et al. (2019), were introduced to compute token-level or document-level similarity using contextualized embeddings . Further improvements came from task-specific metrics fine-tuned on human ratings, such as COMET Rei et al. (2020), COMET-KIWI Rei et al. (2022), BLEURT Sellam et al. (2020), and Prism Thompson and Post (2020). These models learn to directly predict human preferences but often generalize poorly across domains and tasks. General-purpose evaluation has increasingly shifted toward leveraging LLMs as evaluators. Prompt-based approaches such as G-Eval Liu et al. (2023), ChatEval Chan et al. (2024); Hong et al. (2025), and GPT-based judge systems treat LLMs as reference-free annotators that directly assess text quality via carefully crafted prompts. While these methods often achieve high correlation with human judgments, they typically rely on proprietary models like GPT-4, making them costly, and sensitive to prompt phrasing Leiter and Eger (2024). An alternative line of work explores probability-based evaluation metrics, including BARTScore Yuan et al. (2021) and GPTScore Fu et al. (2024). BARTScore and GPTScore share a common formulation: they compute the conditional log-likelihood of a generation given context, using a pretrained language model as the scoring function. However, these evaluation methods inherit the biases and limitations of the underlying model Ohi et al. (2024).
Contrastive Paradigm.
Contrastive methods have recently gained attention in text generation, where the differences between models of varying capacities are leveraged to improve output quality Su et al. (2022); O’Brien and Lewis (2023); Li et al. (2023). Rather than relying on ensemble methods or majority voting, contrastive decoding explicitly uses a weaker model as a contrastive signal to refine or steer the output of a stronger model. This approach has demonstrated success in mitigating common generation issues such as repetition, overconfidence, and incoherence. Moreover, contrastive principles have been extended to model architecture optimization. Shi et al. (2024) propose the Self-Contrast Mixture-of-Experts (SCMoE) framework, where different expert pathways within the same model act as internal contrastive filters, which improves reasoning capabilities and overall model performance. Note that contrastive decoding is also related to language model arithmetic in which arithmetic combinations of LLMs are considered for adjusted generation (Dekoninck et al., 2024). This includes subtraction, which leads to output favored by one LLM but disfavored by the other, which could be leveraged in adversarial settings (Zhang and Eger, 2024).
Inspired by these advancements, as described in §3.3, we extend contrastive principles beyond generation to the domain of evaluation. Instead of refining model outputs, we propose ContrastScore—a novel metric with a different formulation from prior contrastive decoding work.
3 Methodology
The primary goal of this study is to overcome limitations of single-model evaluators, such as capacity bias and over-reliance on model-specific likelihoods, by capturing where strong and weak models disagree in token-level probabilities. This section first reviews generative likelihood-based evaluation and contrastive decoding, then presents the design and formulation of ContrastScore.
3.1 Generative Evaluation
Our evaluation framework builds on probability-based text assessment methods, such as BARTScore, which estimate quality through log-likelihood computations under a pretrained generative model. Given a generated hypothesis h, its quality is evaluated as:
$\text{Score}(\textbf{h}|d,\mathcal{S})=\sum_{t=1}^{m}w_{t}\log P(h_{t}|\textbf{h}_{<t},\mathcal{S},\theta)$ (1)
where $\mathcal{S}$ represents the supplementary text, which may consist of the source text s (in a source-based setting) or the reference output r (in a reference-based setting). $d$ denotes the evaluation dimension (e.g., fluency), $w_{t}$  [^2] Following prior studies Yuan et al. (2021); Qin et al. (2023), $w_{t}$ is weighed equally for all tokens in this work. refers to the token-level weight, and $\theta$ is the model parameters.
3.2 Contrastive Decoding
Assume an expert and amateur model assigning probabilities
$p_{\text{EXP}}^{t}=p(h_{t}\mid h_{<t},\mathcal{S},\theta_{\text{EXP}})$ (2)
$p_{\text{AMA}}^{t}=p(h_{t}\mid h_{<t},\mathcal{S},\theta_{\text{AMA}})$ (3)
to the next token $h_{t}$, where $p_{\text{EXP}}^{t}$ and $p_{\text{AMA}}^{t}$ represent the probabilities of the expert and amateur models, respectively, for $h_{t}$ and where $\theta_{\text{EXP}}$ and $\theta_{\text{AMA}}$ refer to the parameters of the expert and amateur model. Li et al. (2023) propose the contrastive decoding objective
$\text{CD-Score}=\begin{cases}\log{\frac{p_{\text{EXP}}^{t}}{p_{\text{AMA}}^{t}}}&\text{if }h_{t}\in V_{\text{head}}(h_{<t})\\ -\infty&\text{else}\end{cases}$ (4)
Here, $V_{\text{head}}(h_{<t})$ selects the top most likely tokens $x_{t}$ under the expert model given history $h_{<t}$. During next token prediction, CD-Score effectively only considers the most likely tokens of the expert model as candidates, and then assigns modified logits $\log{\frac{p_{\text{EXP}}^{t}}{p_{\text{AMA}}^{t}}}$ to them. The intuition behind CD-Score is to consider the most likely expert tokens as continuation for generation but then choose those tokens among them which the amateur might not favor (e.g., repetitive tokens).
Unfortunately, CD-Score is unsuitable for evaluation because every generated sequence h containing a token $h_{t}$ not in $V_{\text{head}}(h_{<t})$ would receive a score of -$\infty$, essentially being ruled out as a candidate, even though it should have been considered as part of the sequence under evaluation. As an alternative, one might consider the division score $\log{\frac{p_{\text{EXP}}^{t}}{p_{\text{AMA}}^{t}}}$ itself as an objective for evaluation. However, as pointed out by Li et al. (2023), this approach suffers from penalizing many standard text sequences as unlikely, for which both expert and amateur would hold the same probabilities. For example, if both models assign a probability of 0.9 or 0.1 to a token, the score remains the same, making it unable to distinguish between confidently correct and incorrect predictions. As a result, this approach fails to highlight important distinctions between tokens that should matter in evaluation. The division score also becomes instable especially when the amateur probability is close to zero.
3.3 ContrastScore
To address this issue for evaluation, we propose a subtraction-based contrastive formulation with a scaling factor $\gamma\in[0,1]$ that downweights the amateur model. In this way, we naturally retain all tokens in the input sequence for evaluation, as well as a property that favors generations preferred by the expert model, similar to CD-Score. Importantly, the subtraction-based formulation ensures that the scores of individual tokens remain comparable across positions—unlike the division-based approach, where small variations in the amateur model’s likelihood can lead to disproportionately large and unstable scores. By using a small $\gamma$, we control the influence of the amateur model’s probability and mitigate such instability, yielding more reliable evaluation signals. Empirically, we choose a formula for contrastive evaluation that leverages the absolute value distance between the expert and the scaled down amateur:
$\text{ContrastScore}=\sum_{t=1}^{m}w_{t}\log\left(\left|p_{\text{EXP}}^{t}-\gamma p_{\text{AMA}}^{t}\right|\right)$ (5)
At each time step $t$, ContrastScore rewards tokens for which expert and (downweighted) amateur generation probabilities are maximally distinct (as the absolute value is a distance function between two distributions). This would ensure that ContrastScore leverages the strengths of the expert model but removes the limitations of the amateur model; in practice, it might for example lean toward generations that are less repetitive, which the amateur model tends to favor. Note, however, that there are probability ranges where the scaled down amateur would be preferred over the expert—e.g., when $p^{t}_{\text{EXP}}=0$, then our formula would still assign the scaled probability of the amateur to $h_{t}$. With $\gamma=1$, ContrastScore would be indifferent between the expert and the amateur but indiscriminately favor disagreement between them, an undesirable behavior similar to the division score above.
4 Experimental Setup
4.1 Datasets and Tasks
For summarization, we use the SummEval dataset Fabbri et al. (2021), which contains model-generated summaries of CNN/DailyMail articles, as well as QAGS-XSUM Wang et al. (2020), which includes 239 system outputs on the XSUM dataset. For machine translation, we use the MQM22 and MQM23 datasets from the WMT22 Freitag et al. (2022) and WMT23 Freitag et al. (2023) Metrics Shared Tasks. The language pairs involved represent a diverse range of translation scenarios, including high-resource (EN-DE), typologically distant (ZH-EN and EN-RU), and low-resource (HE-EN) settings, thereby enabling a comprehensive evaluation of our approach. See AppendixA.1 for more details.
4.2 Baseline Metrics
We consider the following baseline metrics for comparison: BLEU Papineni et al. (2002), CHRF Papineni et al. (2002), ROUGE 1, ROUGE 2, ROUGE L Lin (2004), COMET Rei et al. (2020)COMET-KIWI Rei et al. (2022), BERTScore Zhang et al. (2020), MoverScore Zhao et al. (2019), and BARTScore Yuan et al. (2021). The details of these metrics can be found in Appendix A.2. In addition, we compare it against Single-model evaluators as well as an Ensemble setting, where the probabilities of the expert and amateur models are averaged per token, using models from the LLaMA and Qwen families to ensure generalizability. The ensemble probability at token $t$, denoted as $p_{\text{Ens}}^{t}$, is computed as:
$p_{\text{Ens}}^{t}=\frac{1}{2}\left(p_{\text{EXP}}^{t}+p_{\text{AMA}}^{t}\right)$ (6)
4.3 Meta-Evaluation
We assess the evaluation metrics in terms of quality, biases, and efficiency.
Correlation with Human Scores (Quality).  The effectiveness of an evaluator is measured using the Pearson correlation between its scores and human scores, where a higher correlation indicates stronger alignment with human judgments. For translation, Pearson correlation assesses the evaluator’s ability to capture differences between candidate translations based on the source texts, comparing them to human judgment scores derived from fine-grained MQM annotations (Freitag et al., 2021). For summarization, we compute Pearson correlation with human judgments across different quality aspects: coherence, fluency, consistency, relevance, and factuality.
Bias Evaluation.  We measure biases in likelihood bias. Recent studies have shown that LLMs tend to favor sentences they deem more likely, often assigning higher evaluation scores and performing better on such sentences, regardless of the specific task or evaluation criteria Ohi et al. (2024); McCoy et al. (2024a, b). Ohi et al. (2024) find that LLM-based evaluators systematically overrate high-likelihood sentences and underrate low-likelihood ones compared to human scores. To quantify this likelihood bias, they propose BiasScore defined as:
$\text{BiasScore}=\rho(\textit{LS},\textit{US})$ (7)
where $\rho$ is the Spearman correlation coefficient, LS (Likelihood Score) represents the probability $P$ assigned by the LLM, and US (Unfairness Score) captures the discrepancy between evaluator scores and human scores. Likelihood bias measures the extent to which an evaluation metric is influenced by the model’s inherent probability estimates. A lower BiasScore indicates that the metric is less dependent on the model’s likelihood biases.
Efficiency.  It is measured by the number of samples processed per second, with a higher processing speed indicating greater efficiency.
4.4 Hyperparameters and Environment
[…]
5 Results
We evaluate ContrastScore in terms of its correlation with human scores, bias, and efficiency, compared to single-model and ensemble methods.
5.1 Correlation with Human Scores
Evaluators QAGS-XSUM SummEval AVG
Factuality Coherence Consistency Fluency Relevance
Baseline ROUGE-1 0.055 0.238 0.079 0.071 0.330 0.154
ROUGE-2 0.121 0.164 0.073 0.066 0.229 0.130
ROUGE-L 0.075 0.207 0.071 0.078 0.303 0.147
MoverScore 0.028 0.129 0.141 0.133 0.280 0.142
BERTScore 0.024 0.342 0.115 0.156 0.372 0.202
BARTScore 0.099 0.450 0.339 0.343 0.428 0.332
LLaMA Single 1B 0.064 0.417 0.549 0.542 0.344 0.383
3B 0.157 0.388 0.575 0.530 0.338 0.398
8B 0.205 0.368 0.588 0.539 0.332 0.406
Ensemble (3B,1B) 0.170 0.427 0.564 0.542 0.344 0.409
(8B,1B) 0.212 0.429 0.581 0.556 0.357 0.427
(8B,3B) 0.241 0.414 0.587 0.544 0.351 0.428
Contrast (3B,1B) 0.262 0.456 0.554 0.551 0.334 0.431
(8B,1B) 0.324 0.461 0.584 0.575 0.368 0.462
(8B,3B) 0.323 0.462 0.587 0.575 0.369 0.463
Qwen Single 0.5B 0.107 0.386 0.536 0.520 0.315 0.373
3B 0.112 0.370 0.544 0.492 0.319 0.367
7B 0.091 0.371 0.560 0.510 0.327 0.372
Ensemble (3B,0.5B) 0.081 0.394 0.568 0.536 0.331 0.382
(7B,0.5B) 0.095 0.386 0.577 0.538 0.333 0.386
(7B,3B) 0.103 0.378 0.584 0.521 0.330 0.383
Contrast (3B,0.5B) 0.128 0.397 0.594 0.546 0.323 0.398
(7B,0.5B) 0.122 0.376 0.608 0.554 0.322 0.396
(7B,3B) 0.134 0.384 0.606 0.557 0.334 0.403
Table 1: Pearson correlation of evaluators with human scores in summarization. boldface represents best overall scores, while underline represents best scores within each model group (Baseline, LLaMA, Qwen). Overall, ContratScore outperforms single and ensemble methods as well as baseline metrics for both LLaMA and Qwen families.
Summarization.  Table 1 presents the Pearson correlation between different evaluators and human scores for summarization. Our results indicate that ContrastScore provides best results overall. Furthermore, ContrastScore consistently surpasses single models while utilizing much fewer parameters. Specifically, ContrastScore with Qwen(3B, 0.5B) outperforms the single Qwen 7B by 7.0%, and ContrastScore with LLaMA(3B, 1B) exceeds single LLaMA 8B by 6.2%. These results reinforce the observation that ContrastScore can efficiently enhance model evaluation by leveraging smaller models. Besides improving overall correlation with human scores, ContrastScore demonstrates notable gains across key summarization dimensions, particularly coherence, fluency, relevance and factuality. When applied to LLaMA(8B, 3B), it enhances coherence by 25.5%, fluency by 6.7%, relevance by 11.1%, and factuality by 57.6%, compared to the single LLaMA 8B. However, its impact on consistency differs across model families.
Machine Translation.  Table 2 presents the Pearson correlation between various evaluators and human scores. The results show that, overall, ContrastScore outperforms both single-model and ensemble approaches. For example, ContrastScore based on Qwen(7B, 3B) improves correlation by 7.1% over the single Qwen 7B model and by 4.6% over the ensemble of Qwen(7B, 0.5B). In contrast, ensembling does not always improve performance: While ensemble methods are generally expected to refine predictions, they can sometimes reduce correlation with human scores. For instance, ensembling LLaMA (8B, 1B) leads to a 3.3% decrease in correlation compared to the LLaMA 8B model. Similar trends can be observed in pairwise accuracy Deutsch et al. (2023), as shown in Table 10 of the Appendix.
Beyond the cases when the amateur model is too weak to provide a meaningful correction signal—such as the low correlation scores observed in the HE-EN pair for both LLaMA 1B and 3B (0.32 and 0.29, respectively), which fail to offer reliable contrastive feedback against LLaMA 8B—our results demonstrate that ContrastScore can achieve correlation scores comparable to, or even exceeding, those of the largest model using only two smaller models. Specifically, ContratScore of Qwen(3B, 0.5B) surpasses single Qwen 7B by 3.1%, improving from 0.449 to 0.463 with only half the parameters.
Evaluators MQM22 MQM23 AVG
EN-DE ZH-EN EN-RU EN-DE ZH-EN HE-EN
Baseline BLEU 0.194 0.156 0.142 0.163 0.085 0.208 0.158
CHRF 0.236 0.156 0.169 0.232 0.063 0.244 0.183
BERTScore 0.263 0.311 0.197 0.325 0.236 0.336 0.278
BARTScore 0.254 0.287 0.201 0.201 0.182 0.317 0.240
COMET 0.476 0.403 0.417 0.432 0.396 0.417 0.424
COMET-KIWI 0.392 0.367 0.354 0.475 0.442 0.395 0.404
LLaMA Single 1B 0.255 0.371 0.286 0.530 0.518 0.320 0.380
3B 0.284 0.366 0.293 0.578 0.526 0.293 0.390
8B 0.363 0.376 0.356 0.632 0.574 0.491 0.465
Ensemble (3B,1B) 0.287 0.371 0.307 0.568 0.538 0.306 0.396
(8B,1B) 0.325 0.382 0.354 0.605 0.575 0.462 0.450
(8B,3B) 0.337 0.379 0.348 0.619 0.574 0.453 0.452
Contrast (3B,1B) 0.338 0.382 0.347 0.599 0.562 0.284 0.419
(8B,1B) 0.392 0.391 0.406 0.641 0.590 0.484 0.484
(8B,3B) 0.383 0.393 0.409 0.639 0.595 0.482 0.483
Qwen Single 0.5B 0.222 0.394 0.294 0.487 0.557 0.347 0.383
3B 0.306 0.413 0.299 0.573 0.594 0.415 0.433
7B 0.326 0.419 0.330 0.600 0.574 0.445 0.449
Ensemble (3B,0.5B) 0.290 0.408 0.315 0.548 0.599 0.421 0.430
(7B,0.5B) 0.301 0.412 0.331 0.567 0.598 0.446 0.443
(7B,3B) 0.333 0.424 0.341 0.599 0.608 0.456 0.460
Contrast (3B,0.5B) 0.342 0.432 0.351 0.590 0.628 0.431 0.463
(7B,0.5B) 0.359 0.435 0.382 0.611 0.628 0.465 0.480
(7B,3B) 0.362 0.439 0.384 0.605 0.629 0.466 0.481
Table 2: Pearson correlation of evaluators with human scores in machine translation. bold represents best overall scores, while underline represents best scores within each model group (Baseline, LLaMA, Qwen). Overall, ContrastScore outperforms single and ensemble methods, as well as baseline metrics for both LLaMA and Qwen families.
5.2 Bias Analysis
Table 3 presents the likelihood bias scores across different evaluators of the LLaMA family. The results clearly demonstrate that ContrastScore substantially mitigates likelihood bias compared to both single-model and ensemble approaches. Specifically, ContrastScore achieves lower likelihood bias than any of the individual models employed. For instance, using LLaMA (8B, 3B), it reduces likelihood bias by 18.0% and 64.3% compared to the single LLaMA 8B model on QAGS-XSUM and MQM22, respectively.
ContrastScore is also effective in mitigating likelihood bias for the Qwen model family, although the improvements are somewhat smaller than those observed with LLaMA. Notably, ContrastScore consistently outperforms all individual Qwen models in reducing likelihood bias, with the exception of the 0.5B model on MQM23. These findings suggest that ContrastScore is highly effective in diminishing the reliance on the inherent likelihood of individual models. See also Tables 8 and 9 in Appendix for more details and breakdown results.
Settings Machine Translation Summarization
MQM22 MQM23 Q-XSUM SummEval
Single
1B 0.342 0.212 0.382 0.348
3B 0.323 0.245 0.289 0.385
8B 0.297 0.352 0.267 0.381
Ensemble
(3B,1B) 0.215 0.123 0.249 0.308
(8B,1B) 0.180 0.152 0.225 0.326
(8B,3B) 0.222 0.229 0.242 0.359
Contrast
(3B,1B) 0.058 0.026 0.233 0.183
(8B,1B) 0.104 0.134 0.220 0.262
(8B,3B) 0.106 0.137 0.219 0.240
Table 3: Likelihood bias scores for machine translation and summarization tasks across LLaMA model family. The lowest overall bias score is boldfaced.
Settings Machine Translation Summarization
MQM22 MQM23 Q-XSUM SummEval
Single
0.5B 0.341 0.252 0.347 0.376
3B 0.442 0.451 0.349 0.392
7B 0.441 0.463 0.373 0.398
Ensemble
(3B,0.5B) 0.331 0.327 0.282 0.379
(7B,0.5B) 0.286 0.294 0.289 0.393
(7B,3B) 0.358 0.381 0.345 0.369
Contrast
(3B,0.5B) 0.307 0.314 0.294 0.236
(7B,0.5B) 0.287 0.302 0.318 0.296
(7B,3B) 0.272 0.287 0.353 0.329
Table 4: Likelihood bias scores for machine translation and summarization tasks across Qwen model family. The lowest overall bias score is boldfaced.
5.3 Efficiency Analysis
Figure 1: Efficiency and quality of ContrastScore with smaller models compared to single larger model on summarization task.
[…]
6 Further Analysis
Source Text 产品防伪标识在哪里
Tokens: Where is the anti - fe iting logo $\bar{\textbf{Log(P)}}$ Rank
Hypothesis 1
Human Rank: 1
Expert 0.2471 0.7305 0.9922 0.02881 0.9805 0.7969 1.000 0.000457 -0.717 2
Amateur 2.265e-06 0.5625 0.9922 0.000335 1.000 1.000 1.000 0.06592 -1.319 1
Contrast 0.2471 0.6758 0.8945 0.02881 0.8789 0.6953 0.8984 0.006134 -0.605 1
Tokens: Where is the anti - fe iting product
Hypothesis 2
Human Rank:2
Expert 0.2471 0.7305 0.9922 0.02881 0.9805 0.7969 1.000 0.002808 -0.618 1
Amateur 2.265e-06 0.5625 0.9922 0.000335 1.000 1.000 1.000 5.841e-05 -1.701 2
Contrast 0.2471 0.6758 0.8945 0.02881 0.8789 0.6953 0.8984 0.002808 -0.647 2
Tokens: What Is The National Debt Limit
Hypothesis 3
Human Rank:3
Expert 0.005951 0.000168 0.2910 4.268e-05 0.001602 0.004944 -2.668 3
Amateur 0.000572 6.845e-08 0.1543 6.482e-07 3.123e-05 0.000140 -4.294 3
Contrast 0.005951 0.000168 0.2754 4.268e-05 0.001602 0.004944 -2.672 3
Table 5: Case Study: Comparison of ContrastScore with Qwen 3B as an expert and Qwen 0.5B as an amateur model for Chinese-to-English (ZH-EN) on MQM23. $\bar{\textbf{Log(P)}}$ denotes mean log-probability across all target tokens.
Weighted Ensemble.   We investigate whether the averaged ensemble baseline used in our experiment is a sufficiently strong setup, or if a weighted combination of model probabilities could yield further improvements. Specifically, we evaluate a weighted ensemble approach that linearly combines the output probabilities of a larger and a smaller model, allowing us to examine whether tuning the weight parameter provides a meaningful advantage over uniform averaging.
$p_{\text{Ens}}^{t}=\gamma\cdot p_{\text{EXP}}^{t}+(1-\gamma)\cdot p_{\text{AMA}}^{t}$ (8)
where $\gamma\in[0,1]$ is the weight factor that controls the contribution of each model. When $\gamma=0.5$, this formulation corresponds to that reported in our main results. To analyze the impact of $\gamma$, we conduct a parameter sweep on the summarization task using LLaMA (3B, 1B), varying $\gamma$ from 0 to 1 in increments of 0.05. The results, presented in Figure 3, show that the weighted ensemble achieves its highest performance at $\gamma=0.55$, which is nearly identical to the performance at $\gamma=0.5$ (i.e., the averaged ensemble baseline). This indicates that the simple averaged ensemble is already a strong and competitive configuration, and that further weighting does not lead to meaningful improvement.
Case Study.
To provide a qualitative analysis, Table 5 presents a case study comparing ContrastScore with a single expert model (Qwen 3B) and an amateur model (Qwen 0.5B) for Chinese-to-English (ZH-EN) machine translation. It can be observed that the expert model misranks Hypothesis 1 as the second-best translation, whereas human evaluators rank it as the best (Rank 1). This misranking is largely influenced by the expert model assigning a very low probability (0.000457) to “logo”, a key token in the translation. In contrast, the amateur model assigns a significantly higher probability (0.06592) to “logo”, demonstrating greater confidence in its relevance. ContrastScore leverages this probability discrepancy to adjust the probability of “logo”, thereby increasing the overall score of Hypothesis 1 and correctly ranking it as the best translation. This case illustrates how ContrastScore mitigates expert model underestimation of critical tokens, leading to improved evaluation alignment with human judgment.
In the case of Hypothesis 3, all methods (i.e., expert, amateur, and ContrastScore) correctly rank it as the worst translation (Rank 3). This consistency occurs because both the expert and amateur models assign low probabilities across all tokens, indicating poor translation quality. Unlike in Hypothesis 1, where significant probability discrepancies require adjustment, ContrastScore makes minimal modifications since the expert and amateur models are already in agreement in their evaluation. This demonstrates that ContrastScore does not introduce unnecessary changes when the models agree, ensuring that it only refines scores when meaningful corrections are needed.
7 Conclusion
In this paper, we introduce ContrastScore, an evaluation metric that leverages structured disagreement between two language models of differing capacities. Unlike conventional LLM-based evaluation methods that rely solely on a single model’s likelihood estimates, ContrastScore incorporates a weaker auxiliary model as a contrastive signal to adjust probability scores for better alignment with human judgments. Extensive experiments demonstrate that ContrastScore consistently achieves stronger correlation with human evaluations than both single-model and ensemble-based baselines. Furthermore, ContrastScore is computationally-efficient and can effectively mitigate likelihood bias, resulting in a more robust evaluation.
Limitations
This work presents the following potential limitations: (1) Our investigation is limited to the LLaMA and Qwen families of models. While ContrastScore demonstrates effectiveness within these two model families, further evaluation across a broader range of large language model architectures is necessary to establish its generalizability. (2) This work relies on token-level probabilities and therefore cannot combine models from different families due to differences in their subword tokenization. In future work, word-level probability could be explored to enable mixing across model families. (3) The metrics explored in this paper are not competitive to the state-of-the-art (SOTA) metrics developed for MT and summarization, such as MetricX (Juraska et al., 2024) or XCOMET-XXL (Guerreiro et al., 2024), which may use much larger models and/or fine-tuning on human annotations. We do not claim in this paper to beat SOTA metrics, but that contrastive principles, involving two models, are superior to non-contrastive principles for evaluation metric design within one evaluation paradigm (BARTScore, in our case). To design competitive metrics, future work should consider involving (considerably) larger models than we did in this work, beyond 8B parameters.
Acknowledgments
We thank the anonymous reviewers for their constructive feedback and helpful suggestions. Xiao Wang is supported by a studentship funded by the Department of Computer Science, University of Manchester.
Appendix A Datasets, Metrics and Prompts
A.1 Datasets
Machine Translation.
MQM22 and MQM23 datasets are annotated by professional translators from the WMT22 Freitag et al. (2022) and WMT23 Freitag et al. (2023) Metrics Shared Tasks, based on the Multidimensional Quality Metrics (MQM) framework Lommel et al. (2014). MQM22 covers 3 language pairs: English-German (EN-DE), Chinese-English (ZH-EN) and English-Russian (EN-RU), and comprises 1315, 1875, and 1315 segments for language pair EN-DE, ZH-EN, EN-RU respectively, with every segments including 15 system outputs. MQM23 covers 3 language pairs: English-German (EN-DE), Chinese-English (ZH-EN) and Hebrew-English (HE-EN), including 460, 1177, and 820 segments and 12, 15, and 12 systems for EN-DE, ZH-EN, HE-EN respectively.
Summarization.
SummEval Fabbri et al. (2021) contains 1600 model-generated summaries of CNN/DailyMail articles, each annotated by 3 experts and 5 crowd workers. It covers aspects of coherence, consistency, fluency, and relevance. QAGS-XSUM Wang et al. (2020) includes 239 system outputs from a fine-tuned BART on XSUM dataset, focusing on factuality aspect.
A.2 Baseline Metrics
BLEU Papineni et al. (2002) is based on the precision of n-grams between the MT output and its reference weighted by a brevity penalty.
CHRF Popović (2015) uses character n-grams instead of word n-grams to compare the MT output with the reference.
ROUGE  Lin (2004) measures the lexical overlap between the hypothesis and reference. We consider 3 variants ROUGE-1, ROUGE-2, and ROUGE-L.
COMET Rei et al. (2020) is a learnt metric that is fine-tuned to produce evaluation scores for a given translation by comparing its representation to source and reference embeddings.
COMET-KIWI Rei et al. (2022) is a reference-free variant of COMET for machine translation evaluation. It uses a multilingual encoder to take only the source and system output as input and predicts a continuous quality score.
BERTScore Zhang et al. (2020) leverages contextual embeddings from BERT to compare words in candidate and reference sentences using cosine similarity.
MoverScore Zhao et al. (2019) measures semantic similarity between a candidate and a reference by combining contextualized embeddings with Earth Mover’s Distance.
BARTScore Yuan et al. (2021) is a generative metric that uses BART to evaluate the generated text by calculating the probabilities of the tokens.
A.3 Prompts
The prompts for the summarization and machine translation tasks are presented in Table 6.
Tasks Prompts
machine translation Translate the following sentence to {target language}:
[…]
Table 6: Prompts for tasks description
A.4 Additional Details
Artifact Licenses
We use the following artifacts:
  • Qwen 2.5 - Qwen License Agreement
  • LLaMA 3.1 and 3.2 - LLaMA 3.1 and 3.2 Community License Agreement
  • COMET22 - Apache 2.0 License
  • COMET-Kiwi - CC BY NC SA 4.0
  • SummEval - MIT License
  • WMT23 - MIT License
Our use of those artifacts complies with license terms and applicable intended use policies.
PII information in data
We did not specifically check whether the datasets used contain PII information. However, we have noticed that WMT23 authors made an effort to mask PII information with special tags. We do not disseminate any new dataset; therefore, PII protection is out of the scope of our work.
Package Versions and Hardware
We use ‘unbabel-comet‘ version 2.2.0 to run baselines for MT evaluation. Experiments were conducted on the university SLURM computing cluster with various available GPUs. For benchmarking purposes, we used H100 GPUs with 96GB of VRAM.
Appendix B Analysis
B.1 The impact of hyperparameter gamma
We study how sensitive our method is to $\gamma$ in Figure 2.
We test ZH-EN language pair of machine translation task, based on LLaMA family. Figure 2 shows that $\gamma\in[0.08,0.2]$ leads to good performance, robust in different model size settings. Furthermore, $\gamma=0.1$ produces the best performance, ensuring a small, controlled refinement.
Figure 2: Exploration of the impacts of $\gamma$. Testing correlation between evaluator score and human score in ZH-EN language pair on MQM23.
B.2 Weighted ensemble
To explore whether the advantages of ContrastScore can be replicated by a simpler method based on probability combination, we evaluate a weighted ensemble approach that linearly merges the output probabilities of the larger and smaller models. The results are presented in the Figure 3.
Figure 3: Exploration of weighted ensemble parameter $\gamma$ for LLaMA-3.2 (3B, 1B) on summarization. Best quality occurs at $\gamma=0.55$ (star), but remains below ContrastScore using the same models. Horizontal lines show individual model performance.
Appendix C Results
C.1 Likelihood Bias
The detailed likelihood bias scores for every datasets of machine translation and summarization tasks are shown in Table 8 and Table 9. ContrastScore demonstrates consistent effectiveness in reducing likelihood bias across both machine translation and summarization tasks, particularly in challenging language pairs and critical evaluation aspects. In machine translation in Table 8, ContrastScore yields substantial improvements for the LLaMA family across diverse language pairs, especially for low-resource or morphologically complex pairs such as EN-RU and HE-EN. In summarization in Table 9, ContrastScore offers the most pronounced improvements in factuality and coherence, which are often the most difficult dimensions for language models. Overall, ContrastScore effectively mitigates likelihood bias compared to both single-model and ensemble baselines in both the LLaMA and Qwen families.
C.2 Efficiency
We use processing speed as an indicator of efficiency, with detailed results for both machine translation and summarization evaluation tasks presented in Table 7. Processing more samples per second indicates higher evaluation efficiency. These results demonstrate that ContrastScore offers substantial gains in evaluation efficiency, especially when leveraging smaller models.
LLaMA SUM MT Qwen SUM MT
Single
1B 63.29 339.04 0.5B 141.96 552.68
3B 47.61 153.25 3B 48.21 194.22
8B 26.43 72.17 7B 26.43 99.84
Ensemble
(3B,1B) 39.51 110.63 (3B,0.5B) 44.94 154.5
(8B,1B) 23.81 66.8 (7B,0.5B) 26.11 93.92
(8B,3B) 19.81 53.95 (7B,3B) 20.35 71.04
Contrast
(3B,1B) 39.12 109.31 (3B,0.5B) 44.06 153.99
(8B,1B) 23.33 66.11 (7B,0.5B) 25.71 93.41
(8B,3B) 19.05 53.45 (7B,3B) 19.61 70.97
Table 7: Processing Speed for Machine Translation and summarization evaluation tasks. Measured in Samples Per Second with batch size of 16 on single H100 GPU.
Settings MQM22 MQM23
EN-DE ZH-EN EN-RU AVG EN-DE ZH-EN HE-EN AVG
LLaMA Single 1B 0.437 0.312 0.276 0.342 0.142 0.237 0.257 0.212
3B 0.426 0.290 0.253 0.323 0.146 0.284 0.304 0.245
8B 0.367 0.287 0.238 0.297 0.181 0.309 0.565 0.352
Ensemble (3B,1B) 0.324 0.219 0.101 0.215 -0.002 0.169 0.202 0.123
(8B,1B) 0.278 0.189 0.074 0.180 0.006 0.148 0.302 0.152
(8B,3B) 0.298 0.227 0.140 0.222 0.079 0.223 0.385 0.229
Contrast (3B,1B) 0.195 0.143 -0.165 0.058 -0.106 0.096 0.088 0.026
(8B,1B) 0.164 0.131 0.016 0.104 -0.044 0.093 0.353 0.134
(8B,3B) 0.187 0.141 -0.011 0.106 -0.051 0.110 0.350 0.137
Qwen Single 0.5B 0.493 0.233 0.296 0.341 0.184 0.214 0.358 0.252
3B 0.580 0.367 0.379 0.442 0.279 0.476 0.598 0.451
7B 0.559 0.383 0.383 0.441 0.333 0.438 0.616 0.463
Ensemble (3B,0.5B) 0.463 0.275 0.256 0.331 0.115 0.370 0.498 0.327
(7B,0.5B) 0.421 0.236 0.202 0.286 0.095 0.304 0.483 0.294
(7B,3B) 0.490 0.303 0.280 0.358 0.206 0.376 0.562 0.381
Contrast (3B,0.5B) 0.432 0.257 0.233 0.307 0.087 0.350 0.505 0.314
(7B,0.5B) 0.415 0.239 0.206 0.287 0.092 0.296 0.518 0.302
(7B,3B) 0.395 0.235 0.187 0.272 0.091 0.292 0.499 0.287
Table 8: Likelihood bias of machine translation task for the LLaMA and Qwen family. ContrastScore can effectively mitigate the likelihood bias compared to both single and ensemble methods on MQM22 and MQM23 datasets.
Settings Q-XUM SummEval
Factuality Coherence Consistency Fluency Relevance AVG
LLaMA Single 1B 0.382 0.104 0.471 0.539 0.279 0.348
3B 0.289 0.169 0.477 0.573 0.321 0.385
8B 0.267 0.161 0.477 0.575 0.311 0.381
Ensemble (3B,1B) 0.249 0.095 0.406 0.488 0.240 0.308
(8B,1B) 0.225 0.119 0.417 0.507 0.261 0.326
(8B,3B) 0.242 0.152 0.450 0.538 0.295 0.359
Contrast (3B,1B) 0.233 -0.008 0.279 0.339 0.122 0.183
(8B,1B) 0.220 0.074 0.337 0.424 0.214 0.262
(8B,3B) 0.219 0.048 0.326 0.400 0.185 0.240
Qwen Single 0.5B 0.347 0.139 0.489 0.567 0.308 0.376
3B 0.349 0.173 0.492 0.570 0.332 0.392
7B 0.373 0.185 0.489 0.584 0.334 0.398
Ensemble (3B,0.5B) 0.282 0.170 0.461 0.567 0.318 0.379
(7B,0.5B) 0.289 0.194 0.464 0.582 0.332 0.393
(7B,3B) 0.345 0.151 0.472 0.560 0.295 0.369
Contrast (3B,0.5B) 0.294 0.017 0.364 0.410 0.152 0.236
(7B,0.5B) 0.318 0.077 0.411 0.488 0.208 0.296
(7B,3B) 0.353 0.112 0.430 0.526 0.249 0.329
Table 9: Likelihood bias of summarization task for the LLaMA and Qwen family. ContrastScore can effectively mitigate the likelihood bias compared to both single and ensemble methods on QAGS-XSUM and SummEval datasets.
Evaluators MQM22 MQM23 AVG
EN-DE ZH-EN EN-RU EN-DE ZH-EN HE-EN
LLaMA Single 1B 0.407 0.493 0.495 0.655 0.625 0.440 0.519
3B 0.415 0.497 0.498 0.661 0.629 0.441 0.524
8B 0.439 0.503 0.515 0.676 0.648 0.511 0.549
Ensemble (3B,1B) 0.416 0.498 0.503 0.665 0.633 0.442 0.526
(8B,1B) 0.426 0.504 0.518 0.674 0.646 0.489 0.543
(8B,3B) 0.430 0.504 0.515 0.675 0.647 0.490 0.544
Contrast (3B,1B) 0.435 0.502 0.509 0.674 0.639 0.438 0.533
(8B,1B) 0.449 0.507 0.528 0.682 0.649 0.496 0.552
(8B,3B) 0.445 0.509 0.530 0.683 0.652 0.496 0.553
Qwen Single 0.5B 0.395 0.500 0.505 0.650 0.637 0.451 0.523
3B 0.417 0.513 0.502 0.665 0.659 0.477 0.523
7B 0.427 0.516 0.517 0.665 0.655 0.494 0.539
Ensemble (3B,0.5B) 0.414 0.510 0.511 0.663 0.656 0.479 0.546
(7B,0.5B) 0.418 0.512 0.518 0.665 0.655 0.491 0.539
(7B,3B) 0.425 0.518 0.518 0.671 0.663 0.497 0.543
Contrast (3B,0.5B) 0.428 0.517 0.519 0.675 0.666 0.482 0.549
(7B,0.5B) 0.435 0.519 0.530 0.676 0.666 0.499 0.548
(7B,3B) 0.435 0.521 0.531 0.676 0.666 0.499 0.554
Table 10: Pairwise Accuracy of evaluators with human scores in machine translation. bold represents best overall scores, while underline represents best scores within each model group (Baseline, LLaMA, Qwen). Overall, ContrastScore outperforms single and ensemble methods, as well as baseline metrics for both LLaMA and Qwen families.